# Mechanism and Emergence of Stacked Attention Heads in Multi-Layer Transformers

### v1

@@ Abstract
In this paper, I introduce the retrieval problem, a simple reasoning task that can be solved only by transformers with a minimum number of layers. The task has an adjustable difficulty that can further increase the required number of layers to any arbitrary value. I demonstrate that large language models can solve the task under different prompting formulations without any fine-tuning. To understand how transformers solve the retrieval problem, I train several transformers on a minimal formulation. I find that successful learning occurs only under the presence of an implicit curriculum. I uncover the learned mechanisms by studying the attention maps in the trained transformers. I also study the training process, uncovering that attention heads always emerge in a specific sequence.
@@ 1 Introduction
I answer Q1 positively by introducing the retrieval problem, as well as a close variant that I term the conditional retrieval problem. I answer Q2 positively by demonstrating that large language models can solve both problems without any specific fine-tuning under multiple prompting formulations. This suggests that large language models have learned a complex mechanism formed by multiple stacked attention heads. To elucidate this mechanism (Q3), I train several transformers on a minimal formulation of the retrieval problem. By studying the attention maps in the trained transformers, I uncover multiple possible mechanisms that I term retrieval heads. Regarding the training process (Q4), I find that retrieval heads emerge only under the presence of an implicit curriculum and always in a specific sequence.
2 Related Work
Single-layer transformers.
Perhaps the most well-studied setting for single-layer transformers is the problem of modular addition. Nanda et al. (2023) show that transformers solve modular addition by arranging the embedding vectors in a circular structure and leveraging the attention mechanism to perform trigonometric operations. Zhong et al. (2024) extend this work by uncovering other algorithms and embedding structures. Even the training dynamics are beginning to be understood, with Ding et al. (2024) studying the survival of initial circular representations and myself (Musat, 2024) proposing an effective theory of the training dynamics by modeling the embeddings as a particle system. Quirke et al. (2023) train a single-layer transformer with three attention heads on the problem of n-digit integer addition. They find that transformers break down the multi-digit addition task into parallel, digit-specific streams, using different algorithms for various digit positions.
Two-layer transformers.
Large language models.
3 Problem Definition
@@
The retrieval problem is directly inspired by the induction problem introduced by Olsson et al. (2022). Given an input sequence ${\bm{a}}{\bm{b}}{\ldots}{\bm{a}}$, the induction problem requires the model to predict the token ${\bm{b}}$. I directly extend this formulation by increasing the number of induction steps to $D$. Given an input sequence ${\bm{x}}_{D-1}{\bm{x}}_{D}\ldots\ldots{\ldots}{\bm{x}}_{1}{\bm{x}}_{2}{\ldots}{\bm{x}}_{0}{\bm{x}}_{1}{\ldots}{\bm{x}}_{0}$, the retrieval problem consists in predicting the token ${\bm{x}}_{D}$. By setting $D=1$, we recover exactly the original induction problem.
@@
I propose an even more general variant of the retrieval problem, which I term the conditional retrieval problem, where each retrieval step could depend on multiple previously retrieved values, not just the last one. For example, given the input sequence ${\bm{x}}{\bm{y}}{\bm{z}}{\ldots}{\bm{a}}{\bm{y}}{\ldots}{\bm{a}}{\bm{x}}{\ldots}{\bm{a}}$, predicting the token ${\bm{z}}$ would constitute a conditional retrieval problem. The retrieval steps in the retrieval problem are perfectly linear, while in the conditional retrieval, they form a directed acyclic graph.
@@ 5 Theoretical Analysis of Information Flow
In this section, I provide formal proof that the retrieval problem requires a minimum number of transformer layers depending on the number of retrieval steps $D$. I model the information flow between different positions during self-attention under the following assumption:
Assumption 1. During self-attention, a position can only attend to another position if they already share a piece of positional or token information.
In other words, a position cannot randomly attend to another position. They must already have some shared piece of information. If they do, information flow is possible between the two positions if (1) the attending position puts the shared information in a subspace of the key vector and (2) the attended position puts the same information in the corresponding subspace the query vector. In Figure 3, I visualize the positions that contain shared information before any transformer layers in the case of $D=5$.
@@
Definitions. Let’s consider all relevant input positions ordered by their reachability from the last token following exactly the paths in Figure 3 (e.g., ${\bm{x}}_{0}{\bm{x}}_{0}{\bm{x}}_{1}{\bm{x}}_{1}{\bm{x}}_{2}{\bm{x}}_{2}\ldots$ ). I denote the residual stream of $i$-th position in this sequence after layer $t$ as ${\bm{r}}_{t,i}$ (e.g., ${\bm{r}}_{0,0}$ is the residual stream of the last token before any transformer layers). I denote by ${\bm{p}}_{i}$ the positional encoding for the $i$-th input pair in the same order of reachability. The initial residual stream ${\bm{r}}_{0,i}$ contains only the token embedding and the positional encoding of the $i$-th position. The token embedding ${\bm{x}}_{i}$ is shared by ${\bm{r}}_{0,2i}$ and ${\bm{r}}_{0,2i+1}$, and the positional encoding ${\bm{p}}_{i}$ is shared by ${\bm{r}}_{0,2i-1}$ and ${\bm{r}}_{0,2i}$. Since the distinction between the token embedding and the positional encoding is not relevant for the information flow, we can consider them together by denoting as ${\bm{e}}_{i}$ the piece of information shared by ${\bm{r}}_{0,i}$ and ${\bm{r}}_{0,i+1}$.
We are interested in the minimum number of layers $t$ such that ${\bm{r}}_{t,0}$ could contain the target token ${\bm{x}}_{D}$. I make a further simplification by introducing an additional assumption that ignores network capacity limitations:
Solving this setting will give us a lower bound on the number of layers required to solve the retrieval problem in the case of a limited network capacity.
@@
Proof. We can show this using mathematical induction. The initial residual stream ${\bm{r}}_{0,i}$ contains only the token embedding and the positional encoding, which represent the consecutive pieces of information $\{{\bm{e}}_{i},{\bm{e}}_{i+1}\}$. During self-attention, the existing contiguous sequence of pieces of information in ${\bm{r}}_{t,i}$ will be merged with other contiguous sequences (assumption 2) that share at least one piece of information with ${\bm{r}}_{t,i}$ (assumption 1). Their union in ${\bm{r}}_{t+1,i}$ remains a contiguous sequence.
@@ 6 Minimal Problem Formulation
I create the input sequences by perfectly interleaving the pairs of symbols forming each retrieval chain, followed by the $N$ query symbols. I randomly shuffle the query vectors. I also shuffle the input pairs from different chains within the same retrieval step. Finally, I concatenate each token embedding with a $K$-dimensional rotary positional encoding (Su et al., 2023). Each input sequence will contain $N(2D+1)$ vectors of dimension $2K$.
The output sequences consist of $N$ vectors, one for each query token. I consider two possible formulations: an implicit curriculum (IC) formulation and a non-IC formulation. In the IC formulation, the target vectors have $DK$ dimensions and contain all the tokens forming the retrieval chain concatenated (except the query token ${\bm{x}}_{0}$). In the non-IC formulation, the target vectors are $K$-dimensional and contain only the last token of each retrieval chain, namely ${\bm{x}}_{D}$.
@@ 7 Implicit Curriculum & Number of Layers
My initial experiments suggest that the implicit curriculum (IC) plays a crucial role in the successful learning of the retrieval problem. To better quantify this effect, I conduct two comprehensive sets of experiments, one for each formulation (IC and non-IC). For each formulation, I train 64 transformers with 1 to 8 layers (8 transformers for each number of layers). I plot the final validation loss averaged across all runs with the same number of layers in Figure 4 (left).
@@ 7.1 Training Details
For each formulation and number of layers, I train 8 transformers following the recipe of Radford et al. (2019). Each transformer has 8 attention heads per layer and residual streams of size 128. I train for 10k steps using the Adam optimizer (Kingma, 2014) with a learning rate of $10^{-3}$, decoupled weight decay of $0.1$ (Loshchilov, 2017), a batch size of 512, $2^{20}$ randomly generated training examples, layer normalization (Ba et al., 2016), no dropout, and mean squared error loss. I measure the final validation loss by averaging the validation loss over the last 100 training steps.
@@ 8 Reverse-Engineering the Circuits Learned
To understand the mechanism learned by transformers to solve the retrieval problem, I train three transformers (denoted as A, B, and C) with 12 layers and only one attention head per layer on the retrieval problem with $D=3$. I then manually reverse-engineer the circuits learned by the transformers by studying their attention maps. The uncovered circuits are depicted in Figure 5.
Figure 5: Manually reverse-engineered circuits from three 12-layer transformers trained on the retrieval problem with $D=3$ and IC.
@@ 8.2 Results
I find that transformers A, B, and C achieve a validation mean squared error of less than $0.01$. By studying the attention maps in the trained transformers, I observe that most attention heads do not perform any useful computation. Only a few attention heads are responsible for the information flow. Their behavior is easily interpretable (see Figure 6). I manually reverse-engineer the entire circuits learned by the three transformers, which are depicted in Figure 5. I display all attention maps in Appendix B.
To validate the circuits, I measure the validation error after ablating the attention maps in the following manner. For the attention heads that do not perform any useful computation, I replace the attention weights with either uniform attention or an identity matrix. For the attention heads that are responsible for the information flow, I construct an attention map that is zero everywhere except for the position that I expect the head to attend to, where it is set to one. I find that the mean squared error increases slightly, but remains below $0.05$ for all transformers. By further ablating as uniform the attention of any apparently useful attention heads, the mean squared error increases to $0.17-0.89$. The only exceptions are the first useful layers of each transformer, which always attend to the previous position. After ablating their attention as uniform, the error stays below $0.1$, suggesting that they do not contribute directly to the final output, but rather enable the information flow in the subsequent layers.
Figure 6: Attention maps for three attention heads in transformer A. The symbols have been assigned letters for better readability.
@@ 9 Emergence of Attention Heads during Training
To better understand how the retrieval heads emerge during training, I train a 24-layer transformer (denoted as Transformer D) on the retrieval problem with $D=4$ and IC. I manually reverse-engineer the circuits learned by Transformer D following the same procedure as before. Afterward, I measure the attention during training for each attention path in the reverse-engineered circuit.
@@ 9.2 Results
Figure 8: Left: manually reverse-engineered circuits of Transformer D. Right: the epoch when the average attention goes above $0.5$ for each attention path.
@@
    After 400 epochs of slow learning, an induction head that can retrieve token B emerges abruptly (Reddy, 2023) on layers 14 and 21. This drives down the first partial loss.
@@
10 Discussion
In this section, I briefly discuss two interesting aspects of my work in relation to language models.
Retreival Problems & Natural Language
In section 4, we saw that large language models achieve great performance on the retrieval problem. However, the retrieval task cannot be explicitly present in the training data since it has not been publicly stated before. How can we reconcile these two facts? One possible explanation is that the retrieval problem is implicitly present as a subproblem in many common tasks such as working with relations between persons, tracking the evolution of a concept, solving mathematical and reasoning problems, programming, and many more. Moreover, the great variety of tasks in natural language data, each requiring different types of retrieval, could act as an implicit curriculum. The emergence of the retrieval mechanism in language models could be explained by the need to work with complex relationships between entities, combined with an implicit curriculum.
Consider the following real-world example from the Wikipedia article on llamas Llamas are social animals and live with others as a herd. [ …] A cria (from Spanish for ‘baby’) is the name for a baby llama, alpaca, vicuña, or guanaco. Crias are typically born with all the females of the herd gathering around.” An autoregressive language model trying to predict the second occurrence of the word herd (rather than flock, group, or pack) would need to first retrieve the fact that crias are llamas, and then use it to retrieve the fact that llamas live in herds. This process is very similar to the retrieval problem with $D=2$.
Emergent Abilities in Large Language Models
Another interesting implication of my work is related to the emergent abilities of large language models (Wei et al., 2022). An ability is emergent if it is not present in smaller models but is present in larger models. Understanding emergence is an important direction because it could potentially allow us to predict what abilities future models may have, as well as provide new insights into how to train more capable language models.

### v2

@@ Abstract
In this paper, I introduce the retrieval problem, a simple yet common reasoning task that can be solved only by transformers with a minimum number of layers, which grows logarithmically with the input size. I empirically show that large language models can solve the task under different prompting formulations without any fine-tuning. To understand how transformers solve the retrieval problem, I train several transformers on a minimal formulation. Successful learning occurs only under the presence of an implicit curriculum. I uncover the learned mechanisms by studying the attention maps in the trained transformers. I also study the training process, uncovering that attention heads always emerge in a specific sequence guided by the implicit curriculum.
@@ 2 My Contribution
I answer Q1 positively by introducing the retrieval problem, as well as a close variant that I term the conditional retrieval problem (Section 3). I answer Q2 positively in Section 4 by empirically showing that large language models can solve both problems without any specific fine-tuning under multiple prompting formulations. In Section 5, I provide a formal proof that the retrieval problem requires a minimum number of transformer layers that is logarithmic in the input size.
@@
This suggests that large language models have learned a complex mechanism formed by multiple stacked attention heads. To elucidate this mechanism (Q3), I train several transformers on a minimal formulation of the retrieval problem (Sections 6 and 7). In Section 8, by studying the attention maps in the trained transformers, I uncover multiple possible mechanisms that I term retrieval heads. Regarding the training process (Q4), I find that retrieval heads emerge only under the presence of an implicit curriculum and always in a specific sequence (Section 9).
The retrieval problem also has important implications for emergent abilities in large language models (Wei et al., 2022), which I discuss in Section 10.
3 The Retrieval Problem
3.1 Definition
@@
The retrieval problem is directly inspired by the induction problem introduced by Olsson et al. (2022). Given a sequence of tokens $\ldots{\bm{a}}{\bm{b}}{\ldots}{\bm{a}}$, the induction problem requires the model to predict the token ${\bm{b}}$. I directly extend this formulation by increasing the number of induction steps to $D$. Given an input sequence $\ldots{\bm{x}}_{D-1}{\bm{x}}_{D}\ldots\ldots{\ldots}{\bm{x}}_{1}{\bm{x}}_{2}{\ldots}{\bm{x}}_{0}{\bm{x}}_{1}{\ldots}{\bm{x}}_{0}$, the retrieval problem consists in predicting the token ${\bm{x}}_{D}$. By setting $D=1$, we recover exactly the original induction problem. Throughout this paper, I also refer to the tokens in the retrieval chain using capital letters (i.e., A for ${\bm{x}}_{0}$, B for ${\bm{x}}_{1}$, C for ${\bm{x}}_{2}$, and so on).
@@
I also propose a more general variant of the retrieval problem, which I term the conditional retrieval problem, where each retrieval step could depend on multiple previously retrieved values, not just the last one. For example, given the input sequence $\ldots{\bm{x}}{\bm{y}}{\bm{z}}{\ldots}{\bm{a}}{\bm{y}}{\ldots}{\bm{a}}{\bm{x}}{\ldots}{\bm{a}}$, predicting the token ${\bm{z}}$ would constitute a conditional retrieval problem. The retrieval steps in the retrieval problem are perfectly linear, while in the conditional retrieval, they form a directed acyclic graph.
3.2 Motivation
The retrieval problem is implicitly present as a subproblem in many common language tasks such as working with relations between persons, tracking the evolution of a concept, solving mathematical and reasoning problems, programming, and many more. Consider the following real-world example from the Wikipedia article on llamas:
Llamas are social animals and live with others as a herd. [ …] A cria (from Spanish for ‘baby’) is the name for a baby llama, alpaca, vicuña, or guanaco. Crias are typically born with all the females of the herd gathering around.”
An autoregressive language model trying to predict the second occurrence of the word herd (rather than flock, group, or pack) would need to first retrieve the fact that crias are llamas, and then use it to retrieve the fact that llamas live in herds. This process is essentially a retrieval problem with $D=2$.
From one point of view, the retrieval problem is essentially about working with relations between entities, which is fundamental for language and reasoning. This makes it an ideal testbed for studying the inner workings of large language models.
@@ 5 Theoretical Analysis of Information Flow
In this section, I provide a formal proof that the retrieval problem requires a minimum number of transformer layers depending on the number of retrieval steps $D$. I model the information flow between different positions during self-attention under the following simplifying assumptions:
Assumption 1. During self-attention, a position can only attend to another position if they already share a piece of information.
This assumption is motivated by the fact that a position can only attend to another position if their key and query vectors align. Constructing aligned key-query pairs is only possible if the two positions already share some information. More precisely, the shared information must be located in the row spaces of the query and key matrices of the attending and attended positions, respectively.
This assumption simplifies the analysis by ignoring network capacity limitations. Solving this setting will give us a lower bound on the number of layers required to solve the retrieval problem in the case of a limited network capacity.
@@
Definition 1. Let’s consider all relevant input positions ordered by their reachability from the last token following exactly the paths in Figure 3 (e.g., ${\bm{x}}_{0}{\bm{x}}_{0}{\bm{x}}_{1}{\bm{x}}_{1}{\bm{x}}_{2}{\bm{x}}_{2}\ldots$ ). Let’s denote this sequence of positions as ${\bm{s}}_{i}$ for $i\in\{0,1,\ldots,2D\}$. In other words, ${\bm{s}}_{2k}$ and ${\bm{s}}_{2k+1}$ are the positions of the second and first occurrences of ${\bm{x}}_{k}$, respectively. For example, ${\bm{s}}_{0}$ is the last position in the input sequence, ${\bm{s}}_{1}$ is the first occurrence of ${\bm{x}}_{0}$, ${\bm{s}}_{2}$ is the second occurrence of ${\bm{x}}_{1}$, ${\bm{s}}_{2D}$ is the position of the target token ${\bm{x}}_{D}$, and so on.
Definition 2. Let’s denote as ${\bm{r}}_{t,i}$ the residual stream at postition ${\bm{s}}_{i}$ after layer $t$. Before any transformer layers, the residual stream contains only the token embedding and positional enconding:
${\bm{r}}_{0,i}={\bm{x}}_{\lfloor i/2\rfloor}+{\bm{p}}_{\lfloor(i+1)/2\rfloor}\quad\text{for all }i\in\{0,1,\ldots,2D\},$ (1)
[…]
Lemma 1. During self-attention, a position can only attend to another position if they already contain a shared token embedding or positional encoding.
Proof. In the retrieval problem, the token values and pair positions are assigned randomly and independently. Knowledge of a token or position does not provide any information about any other token or position. Therefore, the only way for two positions to share information is if they already contain a shared token embedding or positional encoding.
Note that before any attention layers, the position ${\bm{s}}_{i}$ will contain information shared only with positions ${\bm{s}}_{i-1}$ and ${\bm{s}}_{i+1}$. This is because the token embedding ${\bm{x}}_{i}$ is shared by ${\bm{s}}_{2i}$ and ${\bm{s}}_{2i+1}$, while the positional encoding ${\bm{p}}_{i}$ (for $i$-th input pair) is shared by ${\bm{s}}_{2i-1}$ and ${\bm{s}}_{2i}$.
Definition 3. Let’s denote as ${\bm{e}}_{i}$ the piece of information shared by the positions ${\bm{s}}_{i}$ and ${\bm{s}}_{i+1}$ before any transformer layers. Specifically, we define ${\bm{e}}_{i}$ for all $i\in\{0,\ldots,2D\}$ such that ${\bm{e}}_{2k}={\bm{x}}_{k}$ for $k\in\{0,1,\ldots,D\}$ and ${\bm{e}}_{2k-1}={\bm{p}}_{k}$ for $k\in\{1,\ldots,D\}$.
We are interested in the minimum number of layers $t$ such that ${\bm{r}}_{t,0}$ might contain the target token ${\bm{x}}_{D}$, also denoted as ${\bm{e}}_{2D}$.
@@
Proof. We can show this using mathematical induction. The initial residual stream ${\bm{r}}_{0,i}$ contains only the token embedding and the positional encoding, which represent the consecutive pieces of information $\{{\bm{e}}_{i},{\bm{e}}_{i+1}\}$. During self-attention, the existing contiguous sequence of pieces of information in ${\bm{r}}_{t,i}$ will be merged with other contiguous sequences (assumption 2) that share at least one piece of information with ${\bm{r}}_{t,i}$ (lemma 1). Their union in ${\bm{r}}_{t+1,i}$ remains a contiguous sequence.
@@ 6 Minimal Problem Formulation
I create the input sequences by perfectly interleaving the pairs of symbols forming each retrieval chain, followed by the $N$ query symbols. I randomly shuffle the query vectors. I also shuffle the input pairs from different chains within the same retrieval step. Finally, I concatenate each token embedding with a $K$-dimensional rotary positional encoding (Su et al., 2023). Each input sequence will contain $N(2D+1)$ vectors of dimension $2K$. The output sequences consist of $N$ vectors, one for each query token.
@@ 7 Implicit Curriculum & Number of Layers
I consider two possible formulations: an implicit curriculum (IC) formulation and a non-IC formulation. In the IC formulation, the target vectors have $DN$ dimensions and contain all the tokens forming each retrieval chain concatenated (except the query token ${\bm{x}}_{0}$). In the non-IC formulation, the target vectors are $K$-dimensional and contain only the last token of each retrieval chain, namely ${\bm{x}}_{D}$.
My initial experiments suggest that the implicit curriculum (IC) plays a crucial role in the successful learning of the retrieval problem. To better quantify this effect, I conduct two comprehensive sets of experiments, one for each formulation (IC and non-IC). For each formulation, I train 64 transformers with 1 to 8 layers (8 transformers for each number of layers). We plot the final validation loss averaged across all runs with the same number of layers in Figure 4 (left).
@@ 7.1 Training Details
For each formulation and number of layers, I train 8 transformers following the recipe of Radford et al. (2019). Each transformer has 8 attention heads per layer and residual streams of size 128. I train for 10k steps using the Adam optimizer (Kingma, 2014) with a learning rate of $10^{-3}$, decoupled weight decay of $0.1$ (Loshchilov, 2017), a batch size of 512, $2^{20}$ randomly generated training examples, layer normalization (Ba et al., 2016), no dropout, and mean squared error loss. We measure the final validation loss by averaging the validation loss over the last 100 training steps.
@@ 8 Reverse-Engineering the Circuits Learned
To understand the mechanism learned by transformers to solve the retrieval problem, I train three transformers (denoted as A, B, and C) with 12 layers and only one attention head per layer on the retrieval problem with $D=3$. I then manually reverse-engineer the circuits learned by the transformers by studying their attention maps. The uncovered circuits are depicted in Figure 5. I describe my reverse-engineering process in detail in Appendix D.
Figure 5: Reverse-engineered circuits from three 12-layer transformers trained on the retrieval problem with $D=3$ and IC.
@@ 8.2 Results
I find that transformers A, B, and C achieve a validation mean squared error of less than $0.01$. By studying the attention maps in the trained transformers, I observe that most attention heads do not perform any useful computation. Only a few attention heads are responsible for the information flow. Their behavior is easily interpretable (see Appendix B).
I manually reverse-engineer the entire circuits learned by the three transformers, which are depicted in Figure 5. I perform extensive validations of the circuits using ablations. My reverse-engineering process is described in detail in Appendix D.
@@ 9 Emergence of Attention Heads during Training
To better understand how the retrieval heads emerge during training, I train a 24-layer transformer (denoted as Transformer D) on the retrieval problem with $D=4$ and IC. I manually reverse-engineer the circuits learned by Transformer D following the same procedure described in Appendix D. Afterward, I measure the attention during training for each attention path in the reverse-engineered circuit.
@@ 9.1 Training Details
Figure 7: Left: Reverse-engineered circuits of Transformer D. Right: the epoch when the average attention goes above $0.5$ for each attention path.
@@ 9.2 Results
    After 450 epochs of slow learning, an induction head that can retrieve token B emerges abruptly on layers 14 and 21. This drives down the first partial loss.
@@
10 Discussion on Emergent Abilities
The retrieval problem has interesting connections to the emergent abilities of large language models (Wei et al., 2022). An ability is emergent if it is not present in smaller models but is present in larger models. Understanding emergence is an important direction because it could potentially allow us to predict what abilities future models may have, as well as provide new insights into how to train more capable language models.
@@ 10.1 Emergent Abilities are Not a Mirage
10.2 The Implicit Curriculum of Natural Language
Is it possible to understand how training on natural language data leads to reasoning abilities? Chan et al. (2022) previously found that natural language has specific data distributional properties that enable emergent in-context learning in transformers. Is it possible to achieve a similarly insightful understanding of the emergence of reasoning abilities in general, not just in-context learning?
The retrieval problem provides a promising avenue for answering this question. In Sections 7 and 9, we saw that retrieval heads can only emerge gradually, one by one, under the presence of an implicit curriculum that provides a sequence of increasingly complex tasks. This suggests that one property of natural language data is incredibly important for the emergence of reasoning abilities: the presence of a very diverse set of tasks with varying levels of difficulty.
11 Related Work
Single-layer transformers.
Perhaps the most well-studied setting for single-layer transformers is the problem of modular addition. Nanda et al. (2023) show that transformers solve modular addition by arranging the embedding vectors in a circular structure and leveraging the attention mechanism to perform trigonometric operations. Zhong et al. (2024) extend this work by uncovering other algorithms and embedding structures. Even the training dynamics are beginning to be understood, with Ding et al. (2024) studying the survival of initial circular representations and my previous work (Musat, 2024) proposing an effective theory of the training dynamics by modeling the embeddings as a particle system. Quirke et al. (2023) train a single-layer transformer with three attention heads on the problem of n-digit integer addition. They find that transformers break down the multi-digit addition task into parallel, digit-specific streams, using different algorithms for various digit positions.
Two-layer transformers.
Large language models.